OpenStudy (anonymous):

Verify trig the identity Verify the identity. (1-sinx)/cosx=cosx/(1+sinx)

OpenStudy (anonymous):

Heres what we start with: \[\frac{ 1-\sin }{ \cos }=\frac{ \cos }{ 1+\sin }\]We could try cross multiplying which will give us the following: \[\cos^2 = (1-\sin)(1+\sin)\] \[\cos^2 = 1-\sin^2\]Do you see what we could do from here?

OpenStudy (anonymous):

Yes thank you! I couldn't get the first step!

OpenStudy (anonymous):

You're welcome, glad I could help.